OpenStudy (anonymous):

are these independent ??? If P(A) = 0.45, P(B) = 0.25, and P(B|A) = 0.45, are A and B independent? yes , no , or cannot determine ??

jimthompson5910 (jim_thompson5910):

What did you get?

jimthompson5910 (jim_thompson5910):

Keep in mind that if two events are independent, then the two events do not alter each others probabilities

OpenStudy (anonymous):

i guessed no ?

jimthompson5910 (jim_thompson5910):

just a guess or did you have any reason behind that?

jimthompson5910 (jim_thompson5910):

it's ok it was just a guess, I'm just curious

OpenStudy (anonymous):

im guessing because it didnt take into account p(b) but yeah it could be a bad guess lol

jimthompson5910 (jim_thompson5910):

notice how it says P(B) = 0.25

jimthompson5910 (jim_thompson5910):

then it says P(B | A) = 0.45

jimthompson5910 (jim_thompson5910):

the first thing says that the probability of event B happening is 0.25, or a 25% chance

jimthompson5910 (jim_thompson5910):

P(B | A) = 0.45 means that given we know event A has happened, the chances of B happening are 0.45 (or a 45% chance) so clearly knowing that event A has happened has altered the probability of event B happening

jimthompson5910 (jim_thompson5910):

So that shows us that A and B are not independent

jimthompson5910 (jim_thompson5910):

if it told us that P(B) = 0.25 and P(B|A) = 0.25, then this would mean that P(B) = P(B|A) and it would tell us that A and B are independent since A didn't change B but that's not the case here

OpenStudy (raden):

alternative : A and B are independent if satisfies : P(A) * P(B) = P(A and B) we knowed that : P(B|A) = P(A and B)/P(A) so, from here we got P(A and B) = P(A) * P(B|A) = 0.45 * 0.45 = 0.2025 just check the multiply of P(A) * P(B) is equal 0.2025 P(A) * P(B) = 0.45 * 0.25 not equal 0.025

OpenStudy (raden):

because P(A) * P(B) not equal P(A and B) so, it is not independent

OpenStudy (anonymous):

thanks guys :)